OpenStudy (anonymous):

Solve. log(x – 1)16 = 4 but the x-1 is slightly below the log instead of ^

OpenStudy (lgbasallote):

\(\LARGE \log_{(x-1)} 6 = 4\) i'll be quiet :S just wanted to type that

OpenStudy (anonymous):

(x-1)^4 = 16

OpenStudy (anonymous):

x-1=2 x=3

OpenStudy (anonymous):

\[ \log_{x – 1} 16 = 4 \implies (x-1)^4=16=2^4\implies x-1=2 \implies x=3\]